OpenStudy (anonymous):

what is 1/root2 + 1/2 ?

OpenStudy (anonymous):

\[1/\sqrt{2}+1/2\]

OpenStudy (anonymous):

how to make Lcm of that ?

Parth (parthkohli):

Try multiplying both the numerator and denominator in \(\dfrac{1}{\sqrt 2}\) by \(\sqrt{2}\)

OpenStudy (anonymous):

@k.m.imran add like this \[\frac{ 1 }{ a } + \frac{ 1 }{ b } = \frac{ b+a }{ ab }\]

OpenStudy (anonymous):

1/sqrt2+2=sqrt2/2+1/2=(sqrt2+1)/2=(1.414+1)/2=2.414/2=1.212 approx.

OpenStudy (anonymous):

following @ParthKohli's steps you'd get. \[ {\sqrt{2}\over2}+{1\over2}= \] this is now a simpler base to proceed